Driver A pulls onto a highway and travels at a rate of 28.00 m/s. Driver B pulls onto the highway 9.75 minutes later. She accelerates from an initial speed of 25 m/s to 35 m/s, over 95 mins. Both drivers are exiting the highway at the same exit, which is a distance of 4.75 km from where they entered. Determine which driver will exit first.
1
Expert's answer
2020-03-06T10:27:31-0500
"s=v(t_a+T)"

"4750=28(t_a+9.75\cdot 60)"

"t_a=-415\ s"

It means that driver A will exit before driver B pulls onto the highway.

Thus, driver A will exit first.